# Design Principles of Zero-Shot Self-Supervised Unknown Emitter Detectors

### v1

@@ V-C Experiments
In the following we outline the experiment setups for the two scenarios, i.e. with all emitters transmitting the same messages or different messages.
V-C1 Same messages (SM) scenario
@@
For the SM scenario, Section III-B, we employ the WiSig dataset described in Section V-A1. We evaluate both 1D CNN and KAN architectures, each producing feature embeddings of size 20, with DMM configured to use 80 clusters. The four-layer 1D CNN model follows the architecture proposed in [o2018over], while the KAN model is implemented with a grid size of 10, input dimensionality of 512, and spline order of 4. For the Auto-Encoding paradigm, we employ a 1D CNN-based decoder, constructed by replacing convolutional layers with deconvolutional layers and max-pooling operations with up-sampling, following the design in [chen2020one]. The KAN-based Auto-Encoder uses a single-layer KAN architecture comprising 20 input nodes and 512 output nodes. In the contrastive learning framework, Gaussian noise with zero mean and a standard deviation of 0.05 is added to input samples as a data augmentation technique.
@@
For evaluation, we employ a cross-validation procedure in which, iteratively, one of the six emitters is designated as the unknown test emitter. In each iteration, the training dataset is composed of signals from the remaining five emitters, collected across all days and using the first $80\%$ of samples from each day. The test dataset includes data from all six emitters, assembled from the remaining $20\%$ of samples per day. Within this test set, the selected emitter is labeled as unknown, while the other five are treated as known.
@@
Both the 1D CNN and KAN architectures are trained for 100 epochs on the training dataset under three learning paradigms: Deep Clustering, Auto-Encoding, and Contrastive Learning with learning rate $10^{-3}$ and Adam optimizer. During training, evaluation metrics are computed on the current test dataset every 15 epochs. For each evaluation step, metrics are averaged across the six cross-validation iterations. The epoch corresponding to the maximal average performance is then selected, and the metrics at this epoch are reported as final results. This approach follows the common practice in unsupervised and open-set learning, where the final epoch may not represent the best model state due to non-monotonic training dynamics [caron2018deep, chen2020simple]. A similar evaluation strategy was previously adopted in [jain2014multi].
@@
Additionally, we evaluate the SVD with feature dimensionality of 20, varying the number of clusters in DMM among $\{10,40,80,120,160,200,240\}$, as well as KAN-based models with feature sizes $\{2,3,5,10,20\}$, using the same cross-validation procedure.
V-C2 Different messages (DM) scenario
@@
For the different messages scenario,Section III-C, we used the ORACLE dataset described in Section V-A2. We test the approaches using two data modalities: raw I/Q and 2D-Constellation (Figure 4 and Eq. 14) with a grid size of $G=60$. The approaches and models for raw I/Q modulation are the same as for the SM scenario (Sec. V-C1). For the 2D-Constellation modality, we evaluate a four-layer 2D CNN, structurally similar to the previously described 1D variant, along with KANs and their SVD-initialized counterparts from Sec. IV-C. For the Auto-Encoder approach, a 2D CNN decoder is utilized, which is also similar to a 1D CNN Decoder. All feature extractors produce embeddings of size 20, and DMM operates with 200 clusters. For KAN-based architectures, the input layer consists of 3600 nodes. In the contrastive learning setup, data augmentations include rotations of 0, $\pi$, $2\pi$, and $3\pi$ radians, as well as additive Gaussian noise with a standard deviation of $0.05$. Each feature extractor is evaluated under three learning paradigms: Deep Clustering, Auto-Encoding, and Contrastive Learning. We also run PCA method with features size 20 across $\{10,40,80,120,160,200,240\}$ numbers of clusters in DMM. For the evaluation, we use a scheme similar to Sec. V-C1. The only difference is that we randomly shuffle emitters beforehand and consequently choose two emitters as unknown to make five folds in total.
TABLE II: SM scenario: WiSig dataset results
@@
TABLE III: DM scenario: ORACLE dataset results with raw I/Q data modality
@@
TABLE IV: DM scenario: ORACLE dataset results with 2D-Constellation data modality. Rows shaded according to ROC-AUC score (green = best, yellow = medium, unshaded = low).
@@ VI Results
In this section, we present the results of evaluating different machine learning approaches, feature learning modules, and data modalities, following the methodology described earlier. The results are organized as a discussion of the design choices illustrated in Fig. 1(a). Table II summarizes the performance of KAN and CNN architectures across three self-supervised learning paradigms. The Approach and FE Module columns specify the training algorithm and feature extractor’s architecture, respectively. The columns ROC-AUC, NMI, and F1 report the corresponding evaluation metrics (in percentages), whereas Params and FLOPs indicate the number of parameters and the number of floating-point operations required to process a single sample. Tables III and IV follow the same organization as Table II, but report results for the ORACLE dataset with different data modalities.
@@ VI-A The impact of data modality
As it can be seen from Tables II,III and IV, raw I/Q data modality shows high performance in the SM scenario and demonstrates severely degraded performance in the DM scenario, while using 2D-Constellation data modality significantly improves the performance in the DM setup due to the time-invariant representation property (Sec. IV-B).
@@
As shown in Table II for the SM scenario, PCA, KAN trained by AE and CNN-1D trained by DC methods achieve the performance of (ROC-AUC = 98.7%, NMI = 56.2%, F1 = 84.8%), (ROC-AUC = 99.3%, NMI = 58.9%, F1 = 88.0%) and (ROC-AUC = 95.4%, NMI = 51.5%, F1 = 76.0%), respectively, which shows high quality of cluster assignments in DMM measured by NMI, and well-shaped clusters reflected by the ROC-AUC and F1 metrics.
@@
The same methods in the DM set-up on raw I/Q data modality shown in Table III perform significantly worse with (ROC-AUC = 42.1%, NMI = 4.9%, F1 = 3.4%), (ROC-AUC = 50.1%, NMI = 2.1%, F1 = 10.2%) and (ROC-AUC = 50.1%, NMI = 2.4%, F1 = 10.1%) for PCA, KAN trained by AE and CNN-1D trained by DC methods, respectively. This can be attributed to the prevalent semantic information in samples, so self-supervised approaches tend to separate messages instead of emitters.
@@
The results in Table IV for using the 2D-Constellation data modality in the DM scenario for the same methods indicate notable performance improvement across the three metrics: (ROC-AUC = 60.1%, NMI = 34.3%, F1 = 17.1%), (ROC-AUC = 70.6%, NMI = 41.1%, F1 = 38.7%) and (ROC-AUC = 65.8%, NMI = 43.3%, F1 = 24.1%). This confirms that the 2D-Constellation data modality reduces the presence of semantic information, allowing self-supervised approaches to better separate emitters.
VI-B The impact of machine learning approach
In Table II one can see that in the SM scenario with WiSig dataset and raw I/Q data modality, AE learning approach achieves the best performance, followed by PCA and CL. However, AE, DC and CL learning approaches evaluated in the DM scenario using raw I/Q data modality do not show significant differences, as shown in Table III, with PCA performing the worst. Finally, Table IV shows that on sparse 2D-Constellation data modality DC approach is the best, followed with significant margins by CL and PCA approaches.
Quantitatively, the results in Table II confirm the superior performance of the AE approach in the SM scenario using raw I/Q data. The KAN-based AE achieves the highest average scores across all metrics, with (ROC-AUC = 99%, NMI = 59%, F1 = 88%), surpassing both PCA (ROC-AUC = 98%, NMI = 56%, F1 = 85%) and the best CL KAN configuration (ROC-AUC = 98%, NMI = 56%, F1 = 84%). These results indicate that AE captures the signal structure most effectively when message content remains consistent, while PCA provides a strong linear baseline and CL lags slightly behind in feature separability.
In contrast, Table III demonstrates that for the DM scenario under raw I/Q data modality, the overall performances of AE, DC, and CL approaches are nearly indistinguishable, with ROC-AUC values clustered around $50\%$, NMI between 0.5% and 10%, and F1 scores between $9\%$ and $12\%$. PCA yields clearly inferior results (ROC-AUC = 42%, NMI = 5, F1 = 3%), confirming its limited capacity for disentangling features when the message-dependent dynamics dominate the emitter signatures.
Finally, Table IV shows that for the sparse 2D-Constellation data modality, the DC approach achieves the best overall results, particularly for KAN models initialized with SVD weights, reaching (ROC-AUC = 71%, NMI = 41%, F1 = 39%). CL ranks second with (ROC-AUC = 66%, NMI = 34%, F1 = 30%), while PCA, although simpler, remains competitive at (ROC-AUC = 60%, NMI = 34%, F1 = 17%).
Figure 7: KANs vs baseline CNN performance on WiSig Dataset. Blue lines represent average value, blue ticks represent error bars and dashed gray lines show the baseline CNN performance.
VI-C The impact of feature learning module
@@
Table II shows that in the SM scenario, the KAN architecture consistently outperforms CNN-1D across all learning paradigms (AE, DC, and CL). PCA achieves comparable performance while requiring nearly an order of magnitude fewer parameters and FLOPs, highlighting its efficiency as a lightweight baseline. In the DM scenario using the raw I/Q data modality (Table III), KAN-based models exhibit weaker cluster assignment quality compared to CNN-based counterparts, whereas PCA underperforms both deep learning approaches across all metrics. For the 2D-Constellation data modality (Table IV), KAN initialized with SVD weights and trained via DC achieves the best performance, substantially surpassing other architectures. CNN-2D models exhibit stable behavior with SVD initialization when trained using CL. PCA continues to provide a strong baseline with competitive clustering quality while maintaining the lowest computational cost. For all the Raw I/Q constellation with 1D CNN models, KANs are several times more demanding in both parameters and FLOPs. while for 2D Constellation data modality KANs are more effective in terms of FLOPs.
@@
Specifically, in the SM scenario (Table II), the KAN-based AE model achieves the best overall performance, reaching (ROC-AUC = 99.3%, NMI = 59%, F1 = 88%), compared to the CNN-1D variant with (ROC-AUC = 99.2%, NMI = 57%, F1 = 88%). A similar trend is observed for DC and CL paradigms, where KAN variants consistently outperform their CNN counterparts by on average 2244 percentage points in ROC-AUC, NMI and F1 metrics. In the same table, PCA achieves (ROC-AUC = 99%, NMI = 56%,F1 = 85%), closely matching deep learning models while using only 1010k parameters and 0.20.2M FLOPs, requiring approximately 883030 times fewer computational resources than CNN or KAN architectures.
@@
In the DM scenario (Table III), the distinction between models diminishes considerably, with all deep approaches performing close to the chance level: the best ROC-AUC does not exceed $(51.3{\pm}1.7){\%}$ (CL CNN-1D), and F1 values remain around $10{\%}$. Here, KAN architectures perform slightly worse in clustering quality (i.e., NMI = 0.5% for DC KAN) compared to CNN models (i.e., NMI = 2.4% for DC CNN-1D), while PCA drops significantly to (ROC-AUC = 42%, NMI = 5%, F1 = 3%), confirming its inability to separate emitters when signal variability is dominated by the message content.
@@
Finally, in the 2D-Constellation data modality (Table IV), KANs initialized with SVD weights and trained by DC demonstrate a clear advantage, achieving (ROC-AUC = 71%, NMI = 41%, F1 = 39%), which surpasses all other architectures by at least 441010 percentage points across metrics. In contrast, the best CNN-2D model trained with CL reaches (ROC-AUC = 66%, NMI = 34%, F1 = 30%) while PCA achieves (ROC-AUC = 60%, NMI = 34%, F1 = 17%), maintaining competitive performance despite minimal complexity. Collectively, these results confirm that KAN architectures benefit substantially from SVD-based initialization under DC, and that PCA remains an efficient and interpretable baseline delivering strong performance relative to its computational footprint.
@@
Figure 7 compares KAN and CNN performance on the WiSig dataset across KAN feature sizes and learning paradigms. In all cases, KANs exhibit a clear monotonic improvement as the feature dimensionality increases, most notably under the Deep Clustering (DC) and Contrastive Learning (CL) schemes, where ROC-AUC rises from approximately $85\%$ to above $97\%$, and NMI improves by nearly $10$ percentage points when moving from $2$ to $20$ features. AE-based KANs display smaller yet consistent gains, maintaining superior low-dimensional performance compared to other paradigms. Across all metrics, KAN models surpass the CNN baseline (dashed gray line) once the feature dimensionality exceeds $5$, confirming that the adaptive spline-based functional representation in KANs scales more effectively with representational capacity than convolutional filters. These results indicate that KANs can leverage higher-dimensional embeddings to improve discriminative clustering and open-set separability, whereas CNNs reach saturation at lower feature sizes due to their limited nonlinearity and fixed receptive-field structure.

### v2

@@ V-C Experiments
In the following, we outline the experiment setups for the two scenarios, i.e., with all emitters transmitting the same messages or different messages.
V-C1 Same Messages Scenario (SMS)
@@
For the SMS scenario described in Sec. III-B, we employ the WiSig dataset described in Sec. V-A1. We evaluate both CNN-1D and KAN architectures, each producing feature embeddings of size 20, with DMM configured to use 80 clusters. The four-layer CNN-1D model follows the architecture proposed in [o2018over], while the KAN model is implemented with a grid size of 10, input dimensionality of 512, and spline order of 4. For the AE paradigm, we employ a 1D CNN-based encoder and decoder. The former is a convenient CNN model with reducing image size and increasing number of channels, and the latter is constructed by replacing convolutional layers with deconvolutional layers and max-pooling operations with up-sampling, following the design in [chen2020one]. The KAN-based AE uses a single-layer KAN architecture comprising 20 input nodes and 512 output nodes. In the Contrastive Learning framework, Gaussian noise with a mean of zero and a standard deviation of 0.05 is added to input samples as a data augmentation technique along with random rotations and amplifying similarity to [hao2023contrastive].
@@
For evaluation, we employ a cross-validation procedure in which, iteratively, one of the six emitters is designated as the unknown test emitter. In each iteration, the training dataset consists of signals from the remaining five emitters, collected across all days and using the first 80% of samples from each day. The test dataset includes data from all six emitters, assembled from the remaining $20\%$ of samples per day. Within this test set, the selected emitter is labeled as unknown, while the other five are treated as known.
@@
Both the CNN-1D and KAN architectures are trained for 100 epochs on the training dataset under three learning paradigms: Deep Clustering, Auto-Encoding, and Contrastive Learning with a learning rate $10^{-3}$ and an Adam optimizer. During training, evaluation metrics are computed on the current test dataset every 15 epochs. For each evaluation step, metrics are averaged across the six cross-validation iterations. The epoch corresponding to the maximal average performance is then selected, and the metrics at this epoch are reported as final results. This approach follows the standard practice in unsupervised and open-set learning, where the final epoch may not represent the best model state due to non-monotonic training dynamics [caron2018deep, chen2020simple]. A similar evaluation strategy was previously adopted in [jain2014multi].
@@
Additionally, we evaluate the PCA with feature dimensionality of 20, varying the number of clusters in DMM among $\{10,40,80,120,160,200,240\}$, as well as KAN-based models with feature sizes $\{2,3,5,10,20\}$, using the same cross-validation procedure.
V-C2 Different Messages Scenario (DMS)
@@
For the DMS scenario described in Sec. III-C, we use the ORACLE dataset described in Sec. V-A2. We perform the experiments using two data modalities: raw I/Q data and 2D constellation data (Figure 4 and Eq. 11) with a grid size of $G=60$. The approaches and models for raw I/Q modulation are the same as for the SMS as explained in Sec. V-C1. For the 2D constellation modality, we evaluate a four-layer 2D CNN, structurally similar to the previously described CNN-1D variant, along with KAN s and their SVD-initialized counterparts from Sec. IV-C. For the AE approach, a 2D CNN decoder is utilized, which is also similar to a CNN-1D Decoder. All feature extractors produce embeddings of size 20, and DMM operates with 200 clusters. For KAN-based architectures, the input layer consists of 3600 nodes. In the contrastive learning setup, data augmentations include rotations of 0, $\pi$, $\frac{\pi}{2}$, and $\frac{3\pi}{2}$ radians, as well as additive Gaussian noise with a standard deviation of $0.05$. Each FLM is evaluated under three learning paradigms: Deep Clustering, Auto-Encoding, and Contrastive Learning. We also run the PCA method with feature size 20 across $\{10,40,80,120,160,200,240\}$ numbers of clusters in DMM. For the evaluation, we use a scheme similar to Sec. V-C1. The only difference is that we randomly shuffle the emitters beforehand and consequently choose two emitters as unknown, making a total of five folds.
TABLE II: SMS: WiSig dataset results with the raw I/Q data modality.
@@
TABLE III: DMS: ORACLE dataset results with the raw I/Q data modality.
@@
TABLE IV: DMS: ORACLE dataset results with the 2D constellation data modality. Rows shaded according to ROC-AUC score (green = best, yellow = medium, unshaded = low).
@@ VI Results
In this section, we analyze aspects of designing the various components of the workflow illustrated in Figure 3. The results are organized as a discussion of the design choices illustrated in Figure 1(b). Table II summarizes the performance of KAN and CNN architectures across three self-supervised learning paradigms. The Approach and FE Module columns specify the ML approach discussed in Sec. IV-D and FLM from Sec. IV-C, respectively. The columns ROC-AUC, NMI, and F1 report the corresponding evaluation metrics (in percentages), whereas Params and FLOPs columns indicate the number of parameters and the number of floating-point operations required to process a single sample. Tables III and IV follow the same organization as Table II, but report results for the ORACLE dataset with different data modalities.
@@ VI-A The impact of data modality
As can be seen from Tables II, III, and IV, raw I/Q data modality shows high performance in the SMS scenario. It demonstrates severely degraded performance in the DMS scenario, whereas using the 2D constellation data modality significantly improves performance in the DMS setup due to the time-invariant representation property described in Sec. IV.
@@
As shown in Table II for the SMS scenario, PCA, KAN trained with the AE approach and CNN-1D trained with the DC approach achieve the performance of (ROC-AUC = 98.7%, NMI = 56.2%, F1 = 84.8%), (ROC-AUC = 99.3%, NMI = 58.9%, F1 = 88.0%) and (ROC-AUC = 95.4%, NMI = 51.5%, F1 = 76.0%), respectively, which shows high quality of cluster assignments in DMM measured by NMI, and well-shaped clusters reflected by the ROC-AUC and F1 metrics.
@@
The same methods in the DMS setup on the raw I/Q data modality shown in Table III perform significantly worse with (ROC-AUC = 42.1%, NMI = 4.9%, F1 = 3.4%), (ROC-AUC = 50.1%, NMI = 2.1%, F1 = 10.2%) and (ROC-AUC = 50.1%, NMI = 2.4%, F1 = 10.1%) for PCA, KAN trained by AE and CNN-1D trained by DC methods, respectively. This can be attributed to the prevalent semantic information in samples, so self-supervised approaches tend to separate messages instead of emitters.
@@
The results in Table IV for using the 2D constellation data modality in the DMS scenario for the same methods indicate notable performance improvement across the three metrics: (ROC-AUC = 60.1%, NMI = 34.3%, F1 = 17.1%), (ROC-AUC = 70.6%, NMI = 41.1%, F1 = 38.7%) and (ROC-AUC = 65.8%, NMI = 43.3%, F1 = 24.1%). This confirms that the 2D constellation data modality reduces the presence of semantic information, allowing self-supervised approaches to better separate emitters. Overall, 2D constellation data modality shows significant improvement in performance under DMS due to the mitigation of different messages difference as described in Sec. IV-A.
VI-B The impact of Feature Learning Module
@@
Table II shows that in the SMS setup, the KAN architecture slightly but consistently outperforms CNN-1D across all learning paradigms (AE, DC, and CL). PCA achieves comparable performance while requiring nearly an order of magnitude fewer parameters and FLOPs, highlighting its efficiency as a lightweight baseline. In the DMS scenario using the raw I/Q data modality, Table III, the KAN-based models exhibit weaker cluster assignment quality compared to the CNN-based counterparts, whereas PCA underperforms both deep learning approaches across all metrics. For the 2D constellation data modality (Table IV), KAN initialized with SVDweights and trained via DC achieves the best performance, substantially surpassing other architectures. CNN-2D models exhibit stable behavior with SVD initialization when trained using CL. PCA continues to provide a strong baseline with competitive clustering quality while maintaining the lowest computational cost. Comparing CNN-1D and KAN models for the raw I/Q data modality, the latter are several times more demanding in both parameters and FLOPs. However, for the 2D constellation data modality, KAN s are more effective in terms of FLOPs.
@@
Specifically, in the SMS scenario from Table II, the KAN-based AE model achieves the best overall performance, reaching (ROC-AUC = 99.3%, NMI = 59%, F1 = 88%), compared to the CNN-1D variant with (ROC-AUC = 99.2%, NMI = 57%, F1 = 88%). A similar trend is observed for DC and CL paradigms, where KAN variants consistently outperform their CNN counterparts by on average 2244 percentage points in ROC-AUC, NMI and F1 metrics. In the same table, PCA achieves (ROC-AUC = 99%, NMI = 56%,F1 = 85%), closely matching deep learning models while using only 1010k parameters and 0.20.2M FLOPs, requiring approximately 883030 times fewer computational resources than CNN or KAN architectures.
@@
In the DMS scenario from Table III, the distinction between models diminishes considerably, with all deep learning approaches performing close to the chance level: the best ROC-AUC does not exceed $(51.3{\pm}1.7){\%}$ (CL CNN-1D ), and F1 values remain around $10{\%}$. Here, KAN architectures perform slightly worse in clustering quality (i.e., NMI = 0.5% for DC KAN) compared to CNN models (i.e., NMI = 2.4% for DC CNN-1D ). At the same time, PCA drops significantly to (ROC-AUC = 42%, NMI = 5%, F1 = 3%), confirming its inability to separate emitters when signal variability is dominated by the message content.
@@
Finally, in the 2D constellation data modality from Table IV, KAN s initialized with SVD weights and trained by DC demonstrate a clear advantage, achieving (ROC-AUC = 71%, NMI = 41%, F1 = 39%), which surpasses all other architectures by at least 441010 percentage points across metrics. In contrast, the best CNN-2D model trained with CL reaches (ROC-AUC = 66%, NMI = 34%, F1 = 30%) while PCA achieves (ROC-AUC = 60%, NMI = 34%, F1 = 17%), maintaining competitive performance despite minimal complexity. Collectively, these results confirm that KAN architectures benefit substantially from SVD-based initialization under DC, and that PCA remains an efficient and interpretable baseline delivering strong performance relative to its computational footprint.
@@
Figure 7 compares KAN and CNN performance on the WiSig dataset across KAN feature sizes and learning paradigms. In all cases, KAN s exhibit a clear monotonic improvement as the feature dimensionality increases, most notably under the Deep Clustering (DC) and Contrastive Learning (CL) schemes, where ROC-AUC rises from approximately $85\%$ to above $97\%$. NMI improves by nearly $10$ percentage points when moving from $2$ to $20$ features. AE-based KAN s display smaller yet consistent gains, maintaining superior low-dimensional performance compared to other paradigms. Across all metrics, KAN models surpass the CNN baseline (dashed gray line) once the feature dimensionality exceeds $5$, confirming that the adaptive spline-based functional representation in KAN s scales more effectively with representational capacity than convolutional filters. These results suggest that KAN s can utilize higher-dimensional embeddings to enhance discriminative clustering and improve open-set separability. In contrast, CNNs reach saturation at lower feature sizes due to their limited nonlinearity and fixed receptive-field structure. Summing up, Kolmogorov-Arnold networks (KANs) are interpretable and have performance comparable to CNNs, a convenient alternative that suffers from ”black-box” design, across SMS and DMS.
Figure 7: KAN s vs baseline CNN performance on WiSig Dataset. Blue lines represent the average value, blue ticks represent the error bars, and dashed gray lines show the baseline CNN performance.
VI-C The impact of the machine learning approach
In Table II, it is evident that in the SMS scenario with the WiSig dataset and raw I/Q data modality, the AE learning approach achieves the best performance, followed by CL. However, AE, DC, and CL learning approaches evaluated in the DMS setup using the raw I/Q data modality do not show significant differences, as shown in Table III, with PCA performing the worst. Finally, Table IV shows that on the sparse 2D constellation data modality, the DC approach is the best.
Quantitatively, the results in Table II confirm the superior performance of the AE approach in the SMS setup using raw I/Q data. The KAN-based AE achieves the highest average scores across all metrics, with (ROC-AUC = 99%, NMI = 59%, F1 = 88%), surpassing the best CL KAN configuration (ROC-AUC = 98%, NMI = 56%, F1 = 84%). These results indicate that AE captures the signal structure most effectively when message content remains consistent. While PCA provides a strong linear baseline, CL lags slightly behind, with (ROC-AUC = 98%, NMI = 56%, F1 = 84%) for CL and (ROC-AUC = 99%, NMI = 56%, F1 = 85%) for PCA.
In contrast, Table III demonstrates that for the DMS scenario under the raw I/Q data modality, the overall performances of AE, DC, and CL approaches are nearly indistinguishable, with ROC-AUC values clustered around $50\%$, NMI between 0.5% and 10%, and F1 scores between $9\%$ and $12\%$.
Finally, Table IV shows that for the sparse 2D constellation data modality, the DC approach achieves the best overall results, particularly for the KAN models initialized with SVD weights, reaching (ROC-AUC = 71%, NMI = 41%, F1 = 39%). CL ranks second with (ROC-AUC = 66%, NMI = 34%, F1 = 30%), while PCA, although simpler, remains competitive at (ROC-AUC = 60%, NMI = 34%, F1 = 17%). Summarizing, for DMS, the best performing configuration is SVD initialized KAN with deep clustering approach and 2D constellation data modality.
Figure 8: Sensitivity of the DMM to the number of clusters. Without loss of generality, we use the PCA baseline with feature size 20.